# Estimating Fold Changes from Partially Observed Outcomes with Applications in Microbial Metagenomics

### v1

@@ 6 Data analysis: Colorectal cancer microbial metaanalysis
Robust Wald confidence intervals for the coefficients on CRC status are shown in Figure 4. We identify 30 taxa for which we reject the null hypothesis of equal log mean cell concentration for CRC patients and (otherwise alike) non-CRC controls at a false discovery rate of 0.1 based on the robust score test. These taxa are denoted by navy-colored confidence intervals. Taxa are ordered along the x-axis by the proportion of participants in which they are detected, highlighting that significantly differentially abundant taxa are identified across the spectrum of detection levels. Taxa our analysis identifies as more enriched than typical among CRC patients relative to similar controls include an unknown Dialister species estimated to have ratio of mean cell concentration 49 times greater than typical (95% CI $12-210$; $p=1.1\times 10^{-6}$), Porphyromonas uenonis (30 times greater; 95% CI $6.2-150$; $p=0.0025$), Fusobacterium nucleatum s. animalis (24 times greater; 95% CI $11-52$; $p=2.2\times 10^{-5}$), and Gemella morbillorum (23 times greater; 95% CI $7.6-69$; $p=0.0016$). Confidence intervals for all effect sizes that are significantly associated with CRC status at FDR cutoff $0.1$ are given in Figure S7.
We identify a number of taxa associated with CRC status that are detected in relatively high proportion in the study populations. For example, we estimate for B. kashiwanohense, which was detected in 35.7% of study participants, that the ratio of mean stool cell concentrations for this taxon comparing CRC patients to alike controls is 8.6 times lower (95% CI 3.6 – 20.9; unadjusted robust score $p=0.0022$) than the typical such ratio among taxa included in our analysis. This is consistent with a possible protective effect of B. kashiwanohense, but we cannot rule out association due to unmeasured confounding (e.g., through diet) or reverse causation (development of CRC induces relative depletion of B. kashiwanohense). We note that B. kashiwanohense was not identified as a taxon negatively associated with CRC status in the analysis of Wirbel et al. (2019).
We compare our results to those of Wirbel et al. (2019), as well as to 5 other popular differential abundance methods in the microbiome literature, in Figure 6. All panels show robust score p-values obtained from our adjusted model on the x-axis, and p-values obtained from comparator methods on the y-axis. We briefly describe details for each method here; please see Code Appendix for a reproducible workflow. (Top left) corncob (Martin et al., 2020) likelihood ratio test p-values with design matrix as for the proposed method and timing of sampling and study as dispersion covariates; (top center) ALDEx2 (Fernandes et al., 2014) with centered log-ratio transform and design matrix as for the proposed method and 512 Monte Carlo iterations; (top right) DESeq2 (Love et al., 2014) likelihood ratio test with design matrix as for the proposed method and “local” estimation of dispersions; (lower left) IFAA (Li et al., 2021) with design matrix as for the proposed method; (lower center) ANCOM-BC2 (Lin and Peddada, 2020) with design matrix as for the proposed method and without prevalence filtering; and (lower right) the results of Wirbel et al. (2019), who report a blocked Wilcoxon test performed on proportion-scale mOTU data, with blocking on study and timing of sampling relative to colonoscopy (five p-values of zero are altered to $10^{-25}$ for visual comparison). While all methods target slightly different estimands, our proposed methodology is most similar to the ALDEx2 and ANCOM-BC2 target estimands, which is consistent with the relatively high sample Spearman correlation of p-values between our method when compared with ALDEx2 (0.53) and ANCOM-BC2 (0.65).
Finally, we contrast our results with those of Wirbel et al. (2019), finding substantial variation in the ordering of p-values between our proposed method and the blocked Wilcoxon test (Spearman correlation $0.44$). Only 45 of the 94 taxa Wirbel et al. (2019) identified as significantly associated with CRC status are among the 94 taxa most significantly associated with CRC status in our analysis, and of the 30 taxa we highlight at FDR level $0.1$, only 10 are among the 30 most highly significant taxa identified by Wirbel et al. (2019). We believe that this lack of concordance is likely due to two major factors. Firstly, our model adjusts for a larger set of potential confounders than do Wirbel et al. (2019), which may account for our discrepant findings. In addition, we estimate effects on the scale of log mean cell concentration, whereas Wirbel et al. (2019) conduct tests on the read proportion scale. Because a perturbation to the abundance of a small number of taxa on the cell concentration scale induces a dense perturbation on the proportion scale, analyses on the proportion scale may highlight taxa that do not vary in a scientifically interesting way on the cell concentration scale. Similarly, analyses performed on proportion data may fail to detect scientifically meaningful effects on the cell concentration scale because of fluctuations in cell concentration among other taxa.
Figure 6: Negative $\log_{10}$ p-values from distinct approaches to differential abundance analyses based on the Wirbel et al. (2019) dataset. The proposed method’s robust score test p-values are shown on the x-axis, and comparator methods on the y-axis. The proposed method produces smaller p-values in 45%, 68%, 52%, 49%, 51% and 41% of cases for corncob, ALDEx2, DESeq2, IFAA, ANCOM-BC2 and blocked Wilcoxon respectively, with rank correlations of 0.45, 0.53, 0.57, 0.14, 0.65 and 0.44. Wirbel et al. (2019) report some p-values equal to zero; to allow graphing on the log scale, we replace these values with $10^{-25}$. In addition to the missing/uncomputable p-values of corncob, DESeq2 and IFAA, two DESeq2 were so small as to make plotting infeasible; these observations correspond to $(p_{\text{radEmu}},p_{\text{DESeq2}})$ tuples of $(10^{-1.8},10^{-31})$ and $(10^{-2.8},10^{-33})$.
7 Discussion
In this paper we present a regression method for inference on means of a nonnegative multivariate outcome $W$ observed after perturbation by unknown sample-specific scaling terms $z$ and category-specific detection effects $\delta$. Our method is motivated by microbiome sequencing experiments, in which outcomes of interest (microbial cell concentration across many species) are typically observed indirectly via sequencing, which is known to be subject to both sample-specific scaling and taxon-specific distortion (McLaren et al., 2019). However, we anticipate that our method may find application in other settings; in particular, we expect that it may be appropriate for analysis of various types of ’omics measurements, where similar scaling and detection effects are likely to distort measurements.
Our method has a number of advantages over existing methods commonly applied to perform differential abundance analyses in microbiome studies. Firstly, it targets a population estimand that is distinguishable from sample- and category-scaling effects and that exists under realistic conditions (positivity of population mean abundances $\mathbb{E}W_{j}$). Notably, we do not require positivity of true sample abundances $W_{ij}$. Secondly, our method does not require data transformations, such as log-ratio transformations or relative abundance transformations. This allows us to use information regarding the precision of observations in estimation. It also avoids the ad-hoc fixes (e.g., adding a small pseudocount to zero observations, restricting attention to nonzero counts, or attempting to infer whether observed zeroes are due to insufficient read depth or reflect true absence of a taxon) that are necessary to apply log-ratio transformations to data containing zeroes. Additionally, we employ a model-robust score test that controls Type 1 error even in small samples with large excess-Poisson dispersion and high degrees of sparsity. Finally, we incorporate an identifiability constraint that may be flexibly specified to suit varying scientific settings. In particular, we do not require a choice of reference taxon. That said, if such a taxon is available, our method is able to take advantage of this information.
An additional consideration in practical applications for this method is measurement error. While our target estimand $\beta$ can be estimated in the presence of systematic measurement error $\delta$, no such guarantees exist for other kinds of measurement error, such as cross-sample contamination. Accordingly, we encourage the use of decontamination techniques (Davis et al., 2018; Clausen and Willis, 2022) prior to use of our proposed method.
Our method admits several possible extensions. In contexts where it is scientifically plausible that the parameter of interest $\beta_{k}$ is sparse, enforcing sparsity via $\ell_{1}$ penalization is an attractive option. However, we note that if most elements of $\beta_{k}$ are truly equal to zero, $\beta_{k}$ is correctly specified under pseudo-Huber constraints (up to negligible differences between medians and pseudo-Huber centers). In addition, while estimation will be valid in the presence of longitudinal or cluster-dependent observations, it remains to extend our inferential methods to this setting. Finally, it may be of interest to extend our method to address settings in which Neyman-Scott-type problems (Neyman and Scott, 1948) are of concern due to the presence of nuisance parameters whose dimension grows with number of observations.
Open-source software implementing the method (including long and short-form documentation) is available at github.com/statdivlab/radEmu. radEmu abbreviates radEmuAbPill, which denotes “using relative abundance data to estimate multiplicative differences in absolute abundances with partially identified log-linear models”.
S2 Pseudo-Huber Centering
@@ 7 Discussion
S3 Algorithms
S3.1 Unconstrained optimization
Algorithm 1 Coordinate Descent for Maximum Likelihood Estimation
@@
Algorithm 2 Scoring Update with Line Search
@@
      • gradient $S^{j}=X^{T}(Y^{j}-\boldsymbol{\mu}^{j}))$ for $\boldsymbol{\mu^{j}}=[\mu_{1}^{j},\dots,\mu_{n}^{j}]^{T}=[\exp(X_{1}\beta^{j}+z_{1}),\dots,\exp(X_{n}\beta^{j}+z_{n})]^{T}$
@@
      Perform line search to find update $\beta^{j}+\alpha d$ satisfying $l(\beta^{j}+\alpha d)\geq l(\beta^{j})+c\alpha d^{T}\nabla l$ (i.e., the Armijo rule) where $l$ is the log likelihood function evaluated in $\beta^{j}$ with $z$ held fixed, $c>0$ is a constant (by default we use $c=10^{-4}$ as suggested by Nocedal and Wright (1999, Ch. 17), $\nabla l$ is the partial derivative of the log likelihood with respect to $\beta^{j}$, and $\alpha$ is initiated at $0.5$ and halved until the Armijo condition is met. After the Armijo condition is met, optionally rescale $\alpha d$ so that $\max_{k}|\alpha d_{k}|\leq\xi$ for some maximum step length $\xi$ before accepting update $\beta^{j}+\alpha d$ to prevent large steps that may lead to ill conditioning. We by default use $\xi=1$.
Algorithm 3 Penalized Likelihood Estimation via Iterative Maximum Likelihood
@@
      Update $\mathbf{Y}^{(t+1)}$ using $\beta^{(t)}$ and $z$ as laid out in Section 3 of main text
@@
S3.3 Additional Details: Efficient approximate Newton method
@@
S4 Type 1 Error & Power: Additional Results
Distribution $J$ $n=10$ $n=50$ $n=250$
Poisson 10 0.061 0.051 0.046
Poisson 50 0.056 0.050 0.051
Poisson 250 0.049 0.056 0.050
ZINB 10 0.028 0.026 0.032
ZINB 50 0.032 0.039 0.043
ZINB 250 0.025 0.043 0.044
@@
Table S1: Empirical type 1 error for the robust score test at the $0.05$ level by simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 10,000 simulations.
Distribution $J$ $n=10$ $n=50$ $n=250$
Poisson 10 0.107 0.045 0.042
Poisson 50 0.159 0.069 0.053
Poisson 250 0.167 0.080 0.057
ZINB 10 0.252 0.058 0.034
ZINB 50 0.293 0.093 0.057
ZINB 250 0.307 0.098 0.059
Table S2: Empirical type 1 error for robust Wald test at the $0.05$ level by simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 10,000 simulations.
Distribution $J$ $n=10$ $n=50$ $n=250$
Poisson 10 0.918 1.000 1.000
Poisson 50 0.718 1.000 1.000
Poisson 250 0.640 1.000 1.000
ZINB 10 0.034 0.138 0.770
ZINB 50 0.040 0.180 0.872
ZINB 250 0.028 0.194 0.908
@@
Table S3: Empirical power for detection of smaller magnitude signal ($\beta_{kj}-g(\beta_{k})=1$) for robust score test at the $0.05$ level by simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 500 simulations.
Distribution $J$ $n=10$ $n=50$ $n=250$
Poisson 10 0.996 1.000 1.000
Poisson 50 0.996 1.000 1.000
Poisson 250 0.992 1.000 1.000
ZINB 10 0.284 0.276 0.814
ZINB 50 0.366 0.380 0.898
ZINB 250 0.410 0.456 0.938
Table S4: Empirical power for detection of smaller magnitude signal ($\beta_{kj}-g(\beta_{k})=1$) for robust Wald test at the $0.05$ level by simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 500 simulations.
Distribution $J$ $n=10$ $n=50$ $n=250$
Poisson 10 0.882 1.000 1.000
Poisson 50 0.986 1.000 1.000
Poisson 250 0.920 1.000 1.000
ZINB 10 0.096 0.990 1.000
ZINB 50 0.386 0.986 1.000
ZINB 250 0.276 0.922 1.000
Table S5: Empirical power for detection of larger magnitude signal ($\beta_{kj}-g(\beta_{k})=5$) for robust score test at the $0.05$ level by simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 500 simulations.
Distribution $J$ $n=10$ $n=50$ $n=250$
Poisson 10 1.000 1.000 1.000
Poisson 50 1.000 1.000 1.000
Poisson 250 1.000 1.000 1.000
ZINB 10 0.920 1.000 1.000
ZINB 50 0.924 1.000 1.000
ZINB 250 0.962 1.000 1.000
Table S6: Empirical power for detection of larger magnitude signal ($\beta_{kj}-g(\beta_{k})=5$) over $500$ simulations for robust Wald test at the $0.05$ level by simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 500 simulations.
S5 Additional figures
Figure S7: Empirical power to reject $\beta_{kj}=g(\beta_{k})$ at the $0.05$ level for zero-inflated Negative Binomial draws. Results are shown for 100 simulations. The primary purpose of these results is to provide investigators with power estimates, that can then be refined further using our simulation code.

### v2

@@ 9 SI: Proofs of identifiability and asymptotic results
[…]
which is a contradiction. Hence $\beta\not\in G_{\beta^{\star}}$ guarantees the impossibility of $\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta^{\star},\mathbf{z}]$ $=\log\mathbb{E}[\mathbf{Y}|\mathbf{X},\beta,\mathbf{z}^{\star}]$. ∎
Assumption 4.
Let $(X,Y,Z)\overset{iid}{\sim}P^{(0)}$ for unknown true data generating mechanism $P^{(0)}$ with bounded support in $\mathbb{R}^{p}\times\mathbb{R}^{J}\times\mathbb{R}.$ Then $P^{(0)}$ satisfies $\log\mathbbm{E}[Y|X,Z;\beta^{(0)}]=Z\mathbf{1}_{J}+{\beta}^{(0)T}X$ for ${\beta}^{(0)}\in\mathbb{R}^{p\times J}$ with $J$-th column equal to $\mathbf{0}_{J}$.
Assumption 5.
$P^{(0)}$ satisfies $\Sigma_{X}:=Cov(X)$ is positive definite.
Lemma 4.
Under Assumptions 4 and 5, $\beta^{(0)}$ is the unique maximizer of the population criterion $M(\beta):=\mathbbm{E}_{P^{(0)}}\hskip 1.42271ptl_{i}^{\text{profile}}(\beta)$.
Proof.
The law of iterated expectation, applying $\log\mathbbm{E}[Y|X,Z;\tilde{\beta}^{(0)}]=Z\mathbf{1}_{J}+\tilde{X}\tilde{\beta}^{(0)}$, and factorizing gives
$\mathbbm{E}$ ${}_{P^{(0)}}\hskip 1.42271ptl_{i}^{\text{profile}}(\beta)=\mathbbm{E}\left[\mathbbm{E}_{Y|X,Z}\hskip 1.42271ptl_{i}^{\text{profile}}(\beta)\right]$
$=\mathbbm{E}_{X,Z}\left[\mathbbm{E}_{Y|X,Z}\hskip 1.42271pt\sum_{j=1}^{J}\left[Y_{j}\log\left(\frac{\exp(1_{\{j<J\}}X^{T}\beta^{j})}{1+\sum_{j^{\prime}=1}^{J-1}\exp(X^{T}\beta^{j^{\prime}})}\right)-\big{|}\big{|}Y\big{|}\big{|}_{1}\frac{\exp(1_{\{j<J\}}X^{T}\beta^{j})}{1+\sum_{j^{\prime}=1}^{J-1}\exp(X^{T}\beta^{j^{\prime}})}\right]\right]$
$=\mathbbm{E}_{X,Z}\Big{[}\sum_{j=1}^{J}\Big{[}\exp\left(1_{\{j<J\}}X^{T}\beta^{(0)j}+Z\right)\log\left(\frac{\exp(1_{\{j<J\}}X^{T}\beta^{j})}{1+\sum_{j^{\prime}=1}^{J-1}\exp(X^{T}\beta^{j^{\prime}})}\right)$
$\hskip 85.35826pt-\left[\sum_{j^{\prime}=1}^{J}\exp(1_{\{j^{\prime}<J\}}X^{T}\beta^{(0)j^{\prime}}+Z)\right]\frac{\exp(1_{\{j<J\}}X^{T}\beta^{j})}{1+\sum_{j^{\prime}=1}^{J-1}\exp(X^{T}\beta^{j^{\prime}})}\Big{]}\Big{]}$
$=\mathbbm{E}_{X,Z}\exp(Z)\left(1+\sum_{j=1}^{J-1}\exp\left(X^{T}\beta^{(0)j}\right)\right)\times$
$\hskip 14.22636pt\sum_{j=1}^{J}\left[\frac{\exp(1_{\{j<J\}}X^{T}\beta^{(0)j})}{1+\sum_{j^{\prime}=1}^{J-1}\exp(X\beta^{(0)j^{\prime}})}\log\left(\frac{\exp(1_{\{j<J\}}X^{T}\beta^{j})}{1+\sum_{j^{\prime}=1}^{J-1}\exp(X^{T}\beta^{j^{\prime}})}\right)-\frac{\exp(1_{\{j<J\}}X^{T}\beta^{j})}{1+\sum_{j^{\prime}=1}^{J-1}\exp(X^{T}\beta^{j^{\prime}})}\right]$
The sum in the interior of the final expression is of form $\mathbf{c}^{T}\log\mathbf{b}-\mathbf{1}^{T}\mathbf{b}$, which is a strictly convex function maximized at $\mathbf{b}=\mathbf{c}$. Therefore, since $\beta^{(0)}$ maximizes the conditional expectation of the penalized log likelihood at every value of $X$ and $Z$, it follows that $\beta^{(0)}$ is a maximizer of $M(\beta)$. Uniqueness of the maximizer follows from the (strict) convexity of $M(\beta)$. ∎
Assumption 6.
${\beta}^{(0)}$ lies in a closed, bounded subset $\mathbf{\Omega}$ of $\mathbb{R}^{p\times J}$
Theorem 5.
[…]
Proof.
[…]
It remains to show that the maximizer of the penalized likelihood satisfies $M_{n}(\hat{\beta}_{n})\geq M_{n}(\beta^{(0)})-o_{p}(1)$. By (9) and determinant identities, we have
$\frac{1}{n}l_{n}^{\text{penalized}}(\beta)$ $=\frac{1}{n}l_{n}^{\text{profile}}(\beta)+\frac{1}{2n}\left[\log\big{|}\frac{1}{n}\tilde{\mathbf{X}}^{T}\mathbf{V}(\beta)\tilde{\mathbf{X}}\big{|}+p(J-1)\log(n)\right],$
and by definition of $\hat{\beta}_{n}$ we have $\frac{1}{n}l_{n}^{\text{penalized}}(\hat{\beta}_{n})\geq\frac{1}{n}l_{n}^{\text{penalized}}(\beta)$ for any $\beta$. Therefore
$M_{n}(\hat{\beta}_{n})$ $\geq M_{n}(\beta^{(0)})-\frac{1}{2n}\left[\log\big{|}\frac{1}{n}\tilde{\mathbf{X}}^{T}\mathbf{V}(\hat{\beta}_{n})\tilde{\mathbf{X}}\big{|}-\log\big{|}\frac{1}{n}\tilde{\mathbf{X}}^{T}\mathbf{V}(\beta^{(0)})\tilde{\mathbf{X}}\big{|}\right].$
Assumptions 2, 3, and the law of large numbers give that $\frac{1}{n}\tilde{\mathbf{X}}^{T}\mathbf{V}(\beta)\tilde{\mathbf{X}}$ has a stable limit in probability, which we denote $I_{\beta}$. Thus $M_{n}(\hat{\beta}_{n})\geq M_{n}(\beta^{(0)})-o(\frac{1}{n})o_{p}(1)$. The result follows. ∎
Theorem 6.
Under Assumptions 46, the MPLE $\sqrt{n}\left(\hat{\tilde{\beta}}^{(n)}-\tilde{\beta}^{(0)}\right)\overset{\mathcal{D}}{\rightarrow}\mathcal{N}(\mathbf{0},\Sigma)$ for positive definite $\Sigma$.
Proof.
[…]
Theorem 7.
[…]
Proof.
Consistency follows from the continuous mapping theorem applying the function $h(\mathbf{b})=b_{j}-g(\mathbf{b})$ for $h:\mathcal{B}\rightarrow\mathbb{R}$ where $\mathcal{B}$ is the set of $J$-dimensional vectors with $J$-th entry equal to zero. Convergence in distribution follows from the delta method. The matrix $H$ can be constructed by considering the derivative of the appropriate generalization of $h$ to vector-valued arguments in $\mathbb{R}^{p(J-1)\times 1}.$
Theorem 8.
Under Assumptions 46 and either $\text{Var}(Y|X,Z;\beta^{(0)})=\phi\times$ $\text{diag}\left\{\mathbbm{E}[Y|X,Z;\beta^{(0)}]\right\}$ or $\text{Var}(Y|X,Z;\beta^{(0)})=\phi\tilde{\mathbf{V}}$ for unknown constant $\phi$, $\hat{\beta}_{n}$ is asymptotically efficient among asymptotically linear estimators of $\beta^{(0)}$.
Proof.
By McCullagh (1983), an estimator is optimal among linear estimators if it solves the quasi-likelihood equations with $\mathbbm{E}[Y]=\mu(\beta)$ and $\text{Var}[Y]=\phi V(\mu),$ assuming only these moment conditions. The Poisson MLE gives the solution under $\text{Var}(Y|X,Z;\beta^{(0)})=\phi\times\text{diag}\left\{\mathbbm{E}[Y|X,Z;\beta^{(0)}]\right\}$ and the Multinomial MLE gives the solution under $\text{Var}(Y|X,Z;\beta^{(0)})$ $=\phi\tilde{\mathbf{V}}$. We have shown that $\hat{\beta}^{\text{Poisson MLE}}=\hat{\beta}^{\text{Multinomial MLE}}$ when both quantities are well defined, and so by Theorem 5 and Slutsky, $\hat{\beta}_{n}-\hat{\beta}^{\text{Multinomial MLE}}=\hat{\beta}_{n}-\hat{\beta}^{\text{Poisson MLE}}=o_{p}(1)$. Furthermore, $\lim_{n}\sqrt{n}\text{Var}(\hat{\beta}_{n})=\lim_{n}\sqrt{n}\text{Var}(\hat{\beta}^{\text{Multinomial MLE}})$. Therefore, $\hat{\beta}_{n}$ attains the smallest asymptotic variance under $\text{Var}(Y|X,Z;\beta^{(0)})=\phi\tilde{\mathbf{V}}$. But since $\hat{\beta}^{\text{Poisson MLE}}=\hat{\beta}^{\text{Multinomial MLE}}$ and $\hat{\beta}^{\text{Poisson MLE}}$ attains the smallest asymptotic variance under $\text{Var}(Y|X,Z;\beta^{(0)})=\phi\times$ $\text{diag}\left\{\mathbbm{E}[Y|X,Z;\beta^{(0)}]\right\}$, $\hat{\beta}_{n}$ must attain this variance bound also. ∎
Figure 7: $\beta$ is only partially identifiable and cannot be estimated on the “absolute” scale without further assumptions (e.g., a reference category that is known to be equal in mean abundance across groups). Full identifiability is established via a constraint function, allowing us to estimate the log-fold differences in true abundances across groups relative to typical differences. The above illustrates the favorable qualities of centering via the pseudo-Huber smoothed median (right) compared to centering by the mean (middle). $\beta_{j}$ for $j=2,3,4,5$ are unlikely to be of scientific interest (relative to $j=1,6$), and the smoothed median constrains these “uninteresting” taxa closer to zero than mean centering. Note that centering by a smoothed median does not require any assumption of sparsity of true effects $\beta$ (left) to detect taxa changing in abundance.
10 SI: Pseudo-Huber Centering
@@
11 SI: Algorithms and additional details regarding estimation
Algorithm 1 Coordinate Descent for Unconstrained Maximum Likelihood Estimation
@@ 11 SI: Algorithms and additional details regarding estimation
Algorithm 2 Scoring Update with Line Search (unconstrained)
@@
      • gradient $S^{j}=X^{T}(Y^{j}-\boldsymbol{\mu}^{j}))$ for $\boldsymbol{\mu^{j}}=[\mu_{1}^{j},\dots,\mu_{n}^{j}]^{T}=[\exp(X_{1}^{T}\beta^{j}+z_{1}),\dots,\exp(X_{n}^{T}\beta^{j}+z_{n})]^{T}$
@@
      Perform line search to find update $\beta^{j}+\alpha d$ satisfying $l(\beta^{j}+\alpha d)\geq l(\beta^{j})+c\alpha d^{T}\nabla l$ (i.e., the Armijo rule) where $l$ is the log likelihood function evaluated in $\beta^{j}$ with $z$ held fixed, $c>0$ is a constant (by default we use $c=10^{-4}$ as suggested by Jorge and Stephen (2006) in Chapter 17), $\nabla l$ is the partial derivative of the log likelihood with respect to $\beta^{j}$, and $\alpha$ is initiated at $0.5$ and halved until the Armijo condition is met. After the Armijo condition is met, optionally rescale $\alpha d$ so that $\max_{k}|\alpha d_{k}|\leq\xi$ for some maximum step length $\xi$ before accepting update $\beta^{j}+\alpha d$ to prevent large steps that may lead to ill conditioning. We by default use $\xi=1$.
Algorithm 3 Penalized Likelihood Estimation via Unconstrained Iterative Maximum Likelihood
@@
      Update $\mathbf{Y}^{(t+1)}$ using $\beta^{(t)}$ and $z$ as described in Section 3 of main text
@@
11.1 Constrained optimization via an efficient approximate Newton method
@@
12 SI: Type 1 Error & Power: Additional Results
Tables 1 and 2 contain empirical type I error rate and power results for the proposed robust score tests and robust Wald tests on simulated data described in Section 6. Table 3 contains a summary of false discoveries from the permutation analysis described in Section 7.
Test Distribution $J$ $n=10$ $n=50$ $n=250$
Robust score test Poisson 10 0.050 0.050 0.062
Robust score test Poisson 50 0.036 0.038 0.050
Robust score test Poisson 250 0.044 0.050 0.046
Robust score test ZINB 10 0.032 0.026 0.028
Robust score test ZINB 50 0.008 0.038 0.048
Robust score test ZINB 250 0.034 0.066 0.042
Robust Wald test Poisson 10 0.096 0.046 0.056
Robust Wald test Poisson 50 0.146 0.050 0.056
Robust Wald test Poisson 250 0.160 0.068 0.048
Robust Wald test ZINB 10 0.242 0.058 0.030
Robust Wald test ZINB 50 0.268 0.100 0.064
Robust Wald test ZINB 250 0.316 0.112 0.058
@@ 12 SI: Type 1 Error & Power: Additional Results
Table 1: Empirical type 1 error at the 5% level for the robust score test and robust Wald test. Results are stratified by test, simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 500 simulations.
Test Signal magnitude Distribution $J$ $n=10$ $n=50$ $n=250$
Robust score test 1 (small) Poisson 10 0.916 1.000 1.000
Robust score test 1 (small) Poisson 50 0.760 1.000 1.000
Robust score test 1 (small) Poisson 250 0.634 0.998 1.000
Robust score test 1 (small) ZINB 10 0.026 0.150 0.802
Robust score test 1 (small) ZINB 50 0.046 0.214 0.882
Robust score test 1 (small) ZINB 250 0.048 0.196 0.870
Robust score test 5 (large) Poisson 10 0.922 1.000 1.000
Robust score test 5 (large) Poisson 50 0.988 1.000 1.000
Robust score test 5 (large) Poisson 250 0.922 1.000 1.000
Robust score test 5 (large) ZINB 10 0.114 0.978 1.000
Robust score test 5 (large) ZINB 50 0.278 0.988 1.000
Robust score test 5 (large) ZINB 250 0.248 0.918 1.000
Robust Wald test 1 (small) Poisson 10 0.990 1.000 1.000
Robust Wald test 1 (small) Poisson 50 0.998 1.000 1.000
Robust Wald test 1 (small) Poisson 250 0.988 1.000 1.000
Robust Wald test 1 (small) ZINB 10 0.304 0.282 0.818
Robust Wald test 1 (small) ZINB 50 0.330 0.408 0.912
Robust Wald test 1 (small) ZINB 250 0.394 0.432 0.902
Robust Wald test 5 (large) Poisson 10 1.000 1.000 1.000
Robust Wald test 5 (large) Poisson 50 1.000 1.000 1.000
Robust Wald test 5 (large) Poisson 250 1.000 1.000 1.000
Robust Wald test 5 (large) ZINB 10 0.926 1.000 1.000
Robust Wald test 5 (large) ZINB 50 0.938 1.000 1.000
Robust Wald test 5 (large) ZINB 250 0.964 1.000 1.000
@@
Table 2: Empirical power at the 5% level for the robust score test and robust Wald test by test, signal magnitude, simulated distribution, number of taxa $J$, and sample size $n$. Results are shown for 500 simulations. The signal magnitude refers to the true value of $\beta_{2j^{*}}-g(\beta_{2})$ when testing the null hypothesis $\beta_{2j^{*}}-g(\beta_{2})=0$.
Method Minimum First quartile Median Third quartile Maximum
ALDEx2 00 00 00 00 00
ANCOM-BC2 $299$ $310$ $316$ $327$ $340$
ANCOM-BC2 with SS $6$ $11$ $14$ $17$ $28$
CLR t-test 00 00 00 00 $1$
DESeq2 00 $2$ $4$ $9$ $14$
IFAA 00 00 00 00 $2$
radEmu score test 00 00 00 00 00
radEmu Wald test $15$ $22$ $27$ $31$ $46$
Table 3: The number of false discoveries of each method determined via a permutation-based approach applied to the data of Wirbel et al. (2019), with 845 total taxa and a false discovery rate of 10%. The data was constructed in order that all discoveries are false. Summaries of the distribution across 20 permutations are reported.